Operations Report – February 11th

rew 205 Operations Report 11 – FEB – 2019

SOL: 1

Name of person filing report: Veronica Trivino

Non-nominal systems: None

Notes on non-nominal systems: N/A

Generator (hours run): Turned on this evening by 1730 (11th Feb), Stopped by this morning 0730 (11th Feb)

Solar— SOC % – Turned off Gen (11th Feb) 21%, Turned on Gen (11th Feb) 70%

Diesel Reading – 25%

Propane Reading – ~ 50%

Ethanol Free Gasoline – 10 gallons

Water (Auxillary tank) – Not in use gallons

Water (Static tank) – around 88%, ~ 485 gallons

Auxillary to Static tank transfer – No

Gallons transferred: Not Applicable

Water in GreenHab – ~ 288.5 gallons

Water (loft) – 12 marked level

Static to Loft Pump used – yes

Water Meter: 01411118 units

Toilet tank emptied: No

Deimos rover used: No, still not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not Applicable

Beginning charge: Not Applicable

Ending charge: Not Applicable

Currently charging: Not Applicable

Opportunity rover used: Used

Hours: 49.2 hours

Beginning charge: 100%

Ending charge: 86%

Currently charging: Yes

Curiosity rover used: Not Used

Hours: 73.6 hours

Beginning charge: 100%

Ending charge: 100%

Currently charging: Yes

Spirit rover used: Used

Hours: 73.2

Beginning charge: 100%

Ending charge: 85%

Currently charging: Yes

Notes on rovers: None

ATV’s Used: None

Reason for use: Not used

Oil Added? N/A

ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar used and why, where? Yes, Off campus

Crew Car used and why, where? Not used

General notes and comments: None

Summary of internet: Nothing to report

Summary of suits and radios: Suit 10 – Fan broken. Email sent to to Scott for repair aid.

Summary of Hab operations: Nothing to report.

Summary of Green Hab operations: Nothing to report

Summary of Science Dome operations: We were operating the autoclave by the handbook that was provided and as the pressure began to rise to the optimal pressure, it began to drop suddenly. We found that there was a leak out of the threads of the pressure valve that attaches the whole valve assembly to the lid, probably because it was not tightened properly. In order to remedy that, the XO got an adjustable wrench and made a tightening quarter turn on the valve, which caused the leak to stop and pressure, began to rise again. However, a minute later, the upper pressure release segment popped off. Upon inspection, we found that the valve sheared where pressure normally releases. It appears that the valve was already weakened as it sheared cleanly. No threads were damaged, so if a replacement valve assembly can be obtained, the issue should be repairable. The minimum sterilization pressure (17 psi) was obtained only momentarily and no higher pressures were experienced.

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report.

Summary of health and safety issues: EVA 2 encountered 2 cases of overheating crew, one light, one more significant. All treated and crew were healthy again within 25 mins. See more in HSO report.

Questions, concerns and requests to Mission Support: None

Copyright © The Mars Society. All rights reserved. | Main Site